if 327.18 g C8H18 are burned what volume of O2 is needed if the pressure is 663.48 torr and the temperature is 70°C
1
Expert's answer
2020-04-08T13:26:05-0400

Answer: 1,158 m3


Explanation:

Reaction:8Н18 + 25О2 16СО2 + 18Н2О


νС8Н18 =m/M = 327,18 g / 114 g/mol = 2,87 mol

Let's find vO2 by proportion:

2 - 2,87

25 - vO2

Herefrom,

vO2= 2,87 • 25 / 2 = 35,875 mol


Next, let's use Ideal gas law pV = vRT

p = 664,48 torr = 88392,486 Pa

T = 70°C = 343,15 K

pV = vRT, herefrom

VO2 = vRT/p = 35,875 mol • 8,314 J/mol•K • 343,15 K / 88392,486 Pa = 1,158 J/Pa = [Pa = J/m3] = 1,158 m3 or 1158 liters


Hopeful, my answer is comprehensive. Have a nice day and be healthy! :)